OpenStudy (anonymous):

4x^2+20x+25 completing the square

jagr2713 (jagr2713):

is there a answer choice

OpenStudy (wolf1728):

4x^2+20x+25 Move the non-x term to the right 4x^2 +20x = -25 Divide equation by coefficient of x^2 x^2 +5x = -6.25 * take the coefficient of X * divide it by 2 * square that number * then add it to both sides of the equation. Coefficient of x = 5 so 5/2 = 2.5 2.5^2 = 6.25 Adding 6.25 to both sides of the equation: x^2 +5x +6.25 = -6.25 + 6.25 x^2 +5x +6.25 = 0 Take the square root of both sides (x +2.5) * (x +2.5) = 0 x = -2.5 x = -2.5 http://1728.org/quadr2.htm